a roller coaster can travel a maximum speed of 31.0m/s at ground level. Determine the minimum speed the coaster would be travelling at if the highest point of the track is 45.0m off the ground.


1
Expert's answer
2022-04-28T13:13:10-0400

"\frac{mv^2}2=mgh+\frac{mu^2}2,"

"u=\sqrt{v^2-2gh}=8.8~\frac ms."